what should be the length of a nichrome wire of resistance 4.5ohm if the length of a similar wire is 60cm and resistance 2.5 phm
1
Expert's answer
2012-06-07T07:46:06-0400
Let's calculate the resistance of a wire per meter:

Rpm = 2.5 ohm / 0.6 m = 25/6 ohm/m.

So, the lenght of a 4.5 ohm wire is

L = 4.5 ohm / (25/6 ohm/m) = 27/25 m = 1.08 m.